The solubility of potassium trioxochlorate (V) at 20 degree Celsius is 8.0moldm3. calculate

  1. the mass which will dissolve in 250g of water to make a saturated solution.
  2. the mass of potassium trioxochlorate (V) is 20g of saturated solution at 20 degree celsius
1
Expert's answer
2021-03-03T02:01:21-0500


a) at 20 degree Celsius 7.3 g will dissolve in100 g of water  ml

250 g=250 ml

7.3=100

x=250

"x=\frac{7.3\times250}{100}=18.25"

b)с=8mol/l

M=122.548

Vs/s=0.02L

"c=\frac{m}{M\times V}"


"m=c\times M\times V=8\times122.548\times0.02=19.61 g"